OpenStudy (anonymous):

A group of boys and girls took a test. Exactly 2/3 of the boys and exactly 3/4 of the girls passed the test. If an equal number of boys and girls passed the test, what fraction of the entire group passed?

OpenStudy (anonymous):

number of girls = g number of boys = b (2/3)b = (3/4)g Solve: ((2/3)b + (3/4)g) /(b+g)

OpenStudy (anonymous):

Substitution should work. Solve the first equation for one variable and substitute in for that variable in the second equation. Let me see your attempt. =)

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

Give it a shot.

OpenStudy (anonymous):

...

OpenStudy (anonymous):

what do you substiturte it for?